### v1

# When Is Degree Enough? Bounds on Degree-Eigenvector Misalignment in Assortative Structured Networks

Abstract
A tight alignment between the degree vector and the leading eigenvector arises naturally in networks with neutral degree mixing and the absence of local structures. Many real-world networks, however, violate both conditions. We derive bounds on the divergence between the degree vector and the eigenvector in networks with degree assortativity and local mesoscopic structures such as communities, core-peripheries, and cycles. Our approach is constructive. We design sufficiently general degree-preserving rewiring algorithms that start from a neutral benchmark and monotonically increase assortativity and the strength of local structures, with each step inducing a perturbation of the adjacency matrix. Using the Stewart–Sun Perturbation Bound, together with explicit spectral-norm control of the rewiring steps, we derive upper bounds on the angle between the eigenvector and the degree vector for modest levels of assortativity and local structures. Our analytical bounds delineate regions of ‘spectral safety’ in which a node’s degree can be used as a reliable measure of its systemic importance in real-world networks.
Keywords: Eigenvector Localization, Steward-Sun Perturbation Bound, Assortativity, Communities, Core-Periphery, Cycles, Degree-Preserving Rewiring.
1 Introduction
Over the last quarter century, many disciplines—including the social and biological sciences—have embraced the idea that a node’s systemic importance depends not only on how many neighbors it has, but also on where it sits in the network as a whole. In such settings, eigenvector-based measures of centrality provide a natural notion of influence or vulnerability. In practice, however, computing eigenvector centrality requires essentially complete knowledge of the network, which is rarely available. By contrast, degree information is often observed or can be estimated reasonably well. Which is why economists, epidemiologists, and other applied researchers frequently use the degree vector as a proxy for the leading eigenvector. This is entirely reasonable when the two are close. Unfortunately, however, the conditions under which degree and eigenvector coincide are quite restrictive. The network must not only exhibit no assortativity, it must also be devoid of all local structures, including communities, core-periphery patterns, and directed cycles. One would be hard-pressed to find a real-world network that meets these conditions. Firm buyer–seller networks, for instance, exhibit a clear core-periphery structure. Networks of disease transmission sit inside rich social graphs with intricate communities and friendship cycles [^1] See Salathé and Jones (2010) and Volz et al. (2011) for evidence on local structures within the social networks of epidemiological transmission. See Chakraborty et al. (2018) and Bacilieri et al. (2025) for empirical evidence on assortativity and local structures in firm buyer-seller networks.. In such circumstances, the leading eigenvector can diverge from the degree profile, with the degree vector becoming a progressively worse proxy as the network departs from the neutral benchmark. Ideally, we would like to know how large this divergence is before using the degree vector as a proxy for the eigenvector. Which is precisely the question we address in this paper. We derive explicit bounds on the divergence between the eigenvector and the degree vector when the network departs from the neutral benchmark by modest amounts of assortativity, community structure, core-periphery pattern, and directed cycles.
Nearly all of what is said in this paper is built upon the Stewart-Sun Perturbation Bound, which tells us how much the eigenvector can move under a perturbation of the adjacency matrix. In its simplest form, the theorem states that if $\mathbf{A}$ is diagonalizable with a simple leading eigenvalue and we form a perturbed matrix $\widetilde{\mathbf{A}}:=\mathbf{A}+\boldsymbol{\Omega}$, then the sine of the angle between the corresponding leading eigenvectors is bounded above by a constant times $\|\boldsymbol{\Omega}\|_{2}$, the spectral norm of the perturbation. Taken at face value, this is a statement about eigenvectors before and after perturbation. It says nothing about degree vectors. Note, however, that under the neutral benchmark, the leading eigenvector of $\mathbf{A}$ is well approximated by the degree vector. Therefore, if we restrict perturbations to degree-preserving rewirings, the Stewart-Sun bound on the distance between the eigenvectors of $\mathbf{A}$ and $\widetilde{\mathbf{A}}$ is equivalent to a bound on the distance between the degree vector and the eigenvector of the perturbed graph $\widetilde{\mathbf{A}}$. We develop this insight by constructing four families of degree-preserving rewiring schemes that progressively introduce (i) assortative mixing, (ii) community structures, (iii) core-periphery patterns, and (iv) directed cycles. For each scheme we explicitly bound $\|\boldsymbol{\Omega}\|_{2}$ in terms of standard scalar measures of these features—Newman’s assortativity coefficient for degree correlations, a modularity-type quantity for communities, a core-periphery contrast parameter, and normalized counts of cycles. Substituting these into the Stewart–Sun inequality yields explicit, interpretable bounds on the divergence between degree and eigenvector centrality as a function of the strength of assortativity and local structure. These are the first analytical results that link commonly used network statistics to quantitative guarantees on the accuracy of degree-based proxies for eigenvector centrality.
We are not the first to study how assortativity-inducing degree-preserving rewiring influences the eigenstructure of a network. Some years ago, Van Mieghem et al. (2010) used degree-preserving rewiring to systematically tune assortativity and examine how the eigen spectrum responds. Our analysis is close in spirit but differs in two key respects. First, we extend the rewiring-based perspective beyond assortativity to a broader class of local structures. Second, and more importantly, we embed these constructions within a perturbation-theoretic framework by linking the resulting perturbation matrices to the Stewart–Sun eigenvector bound. In this sense, our paper bridges several strands of work in network science and spectral graph theory: rewiring-based control of assortativity, the study of eigenvector localization, and the use of spectral perturbation theory to assess the robustness of centrality measures. It is perhaps worth noting that we take a constructive approach to proof, i.e., we work through algorithms that alter a graph towards desired geometries. This allows us to characterize the angle of deviation for finite-sized graphs with specific spectral properties.
The paper is organized as follows. Section 2 sets notation and collects the spectral preliminaries needed for our main arguments, including the Stewart–Sun perturbation bound. Section 3 bounds the moments of the angle of deviation with the moments of network statistics. Section 4 details the assortativity-inducing degree-preserving rewiring process. It then applies the moment-bound proposition to assortativity, while noting how the constants in the bound depend on certain structural features of the network. Section 5 details the degree-preserving rewiring processes used to generate three distinct local structures: communities, core-periphery patterns, and cycles. We generalize each of these structures, with $K$ communities, $k$ cycles, and perhaps most interestingly, fractal core-periphery wherein the periphery itself consists of ‘miniature’ core-periphery structures. We then apply the moment bound proposition to bound the angle of deviation with measures of communities, core-periphery, and cycles. Section 6 discusses how the heavy tails of the degree distribution shape the bounds on the angle of deviation between the degree-vector and eigenvector. Section 7 offers concluding remarks.
2 Spectral Preliminaries and the Stewart–Sun Perturbation Bound
2.1 Basic notation
$\mathbb{E}$ stands for expectation and $\mathbb{V}$ for variance. Matrices are denoted by bold capital letters, vectors by bold lowercase letters, and scalars by plain letters: for example, $\mathbf{X}$ is a matrix, $\mathbf{x}$ a vector, and $x$ a scalar. We write $\mathbf{X}^{\top}$ and $\mathbf{X}^{-1}$ for the transpose and inverse of $\mathbf{X}$, respectively. For vectors $\mathbf{x}\in\mathbb{R}^{n}$ we use $\|\mathbf{x}\|_{2}$ to denote the Euclidean norm. Given a degree sequence $d_{1},d_{2},...,d_{n}]$, let $\mathbf{d}$ denote the $\ell_{2}$-normalized degree vector corresponding to this degree sequence. $\mathbf{d}$ may be interpreted as in-degree or out-degree based on context. For any square matrix $\mathbf{X}$ we write $\lambda_{k}(\mathbf{X})$ for its $k$-th eigenvalue, ordered by nonincreasing absolute value,
$|\lambda_{1}(\mathbf{X})|\;\geq\;|\lambda_{2}(\mathbf{X})|\;\geq\;\cdots$
and when the matrix is clear from context we abbreviate $\lambda_{1}(\mathbf{X})$ to $\lambda_{1}$. Given any square matrix $\mathbf{X}\in\mathbb{R}^{n\times n}$, we define its spectral norm by
$\|\mathbf{X}\|_{2}\;:=\;\sup\bigl\{\|\mathbf{X}\mathbf{z}\|_{2}:\mathbf{z}\in\mathbb{R}^{n},\ \|\mathbf{z}\|_{2}=1\bigr\}$
$\mathbf{v}^{\mathrm{R}}$ and $\mathbf{v}^{\mathrm{L}}$ denote the unit leading right and left eigenvectors of $\mathbf{A}$, respectively. For brevity, we write $\mathbf{v}$ for whichever of these two unit eigenvectors is relevant in a given context. The Euclidean norm of a vector $\mathbf{x}$ is denoted by $\mathbf{x}||_{2}$.
Definition 1 (Distance between two vectors as an angle).
For two nonzero vectors $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$ we define their angular distance as the acute angle between them
$\theta_{\mathbf{x},\mathbf{y}}\;:=\;\arccos\!\biggl(\frac{|\langle\mathbf{x},\mathbf{y}\rangle|}{\|\mathbf{x}\|_{2}\,\|\mathbf{y}\|_{2}}\biggr)\;\in\;[0,\tfrac{\pi}{2}]$
which is the standard notion of angle between vectors in Euclidean space. When $\mathbf{x}$ and $\mathbf{y}$ are unit vectors, this reduces to $\theta_{\mathbf{x},\mathbf{y}}=\arccos(|\langle\mathbf{x},\mathbf{y}\rangle|)$. Whenever we compare or distance between a degree vector and an eigenvector (for example, an alignment $\theta_{\mathbf{d},\mathbf{v}}$) is to be understood as comparing ‘out-degree and right-eigenvector’ or ‘in-degree and left-eigenvector’. $\circ$
2.2 Neutral matrix and its alignment with the degree vector
Definition 2 (Neutral network).
A baseline neutral network is an unweighted directed graph on $n$ nodes with adjacency matrix $\mathbf{A}\in\mathbb{R}_{\geq 0}^{n\times n}$. We assume that $\mathbf{A}$ is irreducible, that the Perron root $\lambda_{1}(\mathbf{A})$ is simple, and that $\mathbf{A}$ is diagonalizable, so that the Perron eigenvectors are well-defined up to scale. [^2] We work throughout in the Perron–Frobenius regime and assume that the rewiring steps we consider are sufficiently few and local that the graph remains irreducible along the trajectory. Equivalently, one may restrict attention to rewiring moves that preserve strong connectivity (e.g., by rejecting swaps that break irreducibility).
Neutrality means that the adjacency matrix contains no systematic structure beyond that implied by its in- and out-degree sequences. More specifically, let $\mathbf{d}^{\mathrm{out}}$ be an out-degree sequence and $\mathbf{d}^{\mathrm{in}}$ be an in-degree sequences. Let $m:=\mathbf{1}^{\top}\mathbf{d}^{\mathrm{out}}$ denote the number of directed edges. We take the neutral baseline to be the degree-only (rank-one) matrix
$\mathbf{A}\;:=\;\frac{\mathbf{d}^{\mathrm{out}}(\mathbf{d}^{\mathrm{in}})^{\top}}{m}$
$\circ$
Proposition 1 (Degree vector as a proxy for the leading eigenvector in neutral networks).
Let $\mathbf{v}$ denote the unit Perron eigenvector of $\mathbf{A}$ , and let $\mathbf{d}$ denote the corresponding unit degree vector. Then
$\theta_{\mathbf{v},\mathbf{d}}\;\approx\;0$
so that degree centrality coincides with eigenvector centrality in the neutral baseline (Newman, 2010, Sec. 7.8). Note that when $\mathbf{v}$ is the right eigenvector, then $\mathbf{d}$ is the out-degree, equivalently for the left eigenvector and the in-degree. $\circ$
2.3 Perturbation matrices and their spectral norms
Definition 3 (Spectral gap).
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be a matrix with simple eigenvalues $\lambda_{1},\dots,\lambda_{n}$, with $\lambda_{1}$ as the leading eigenvalue. The spectral gap of $\mathbf{A}$ is defined as
$\gamma(\mathbf{A})\;:=\;\min_{j\geq 2}\bigl|\lambda_{1}-\lambda_{j}\bigr|$
The spectral gap measures how well separated the leading eigenvalue is from the rest of the spectrum. $\circ$
Definition 4 (Perturbation matrices).
Consider a sequence of perturbation matrices
$\boldsymbol{\Delta}^{(1)},\,\boldsymbol{\Delta}^{(2)},\,\dots,\,\boldsymbol{\Delta}^{(t)}\in\mathbb{R}^{n\times n}$
with entries in $\{-1,0,1\}$. Each $\boldsymbol{\Delta}^{(i)}$ represents a single degree-preserving rewiring step involving two edges and has exactly four nonzero entries: two equal to +1+1 and two equal to 1-1. We write the cumulative perturbation after $t$ steps as
$\boldsymbol{\Omega}^{(t)}\;:=\;\sum_{i=1}^{t}\boldsymbol{\Delta}^{(i)}.$
Accordingly, the adjacency matrix after $t$ perturbations is
$\mathbf{A}^{(t)}\;:=\;\mathbf{A}\;+\;\boldsymbol{\Omega}^{(t)}.$
We assume that $\mathbf{A}^{(t)}$ is irreducible and diagonalizable for all $t$ (equivalently, we reject any rewiring that breaks these properties (since $\mathbf{A}^{(0)}=\mathbf{A}$ is assumed to be irreducible and diagonalizable). $\circ$
Assumption 1 ($r$-bounded participation rewiring).
Each node participates in at most $r$ rewiring steps. Equivalently, there exist permutation matrices $\mathbf{\Pi}_{1},\dots,\mathbf{\Pi}_{r}$ and a partition of $\{1,\dots,k\}$ into $r$ subsets $I_{1},\dots,I_{r}$ such that for each $s\in\{1,\dots,r\}$
$\mathbf{\Pi}_{s}^{\top}\Big(\sum_{i\in I_{s}}\boldsymbol{\Delta}^{(i)}\Big)\mathbf{\Pi}_{s}$
is block-diagonal with one block for each $\boldsymbol{\Delta}^{(i)}$, $i\in I_{s}$. $\circ$
Proposition 2 (Spectral norm bound for the cumulative rewiring perturbation).
If each vertex participates in at most $r\in\mathbb{N}$ swaps, then
$\|\boldsymbol{\Omega}^{(t)}\|_{2}\;\leq\;2r$
Proof.
Fix a vertex $u$ and let $s(u)$ be the number of swaps in which $u$ participates. In any single degree-preserving swap, a participating vertex has exactly two incident edges whose adjacency entries flip (one $1\!\to\!0$ and one $0\!\to\!1$). Consequently, in the cumulative matrix $\boldsymbol{\Omega}^{(t)}$, each time $u$ participates it can contribute at most two nonzero entries of magnitude $1$ in row $u$, and likewise at most two such entries in column $u$. Hence
$\sum_{j}\bigl|\Omega^{(t)}_{uj}\bigr|\;\leq\;2s(u),\qquad\sum_{i}\bigl|\Omega^{(t)}_{iu}\bigr|\;\leq\;2s(u)$
Let $s_{\max}:=\max_{u}s(u)\leq r$. Taking maxima over $u$ gives the induced $\ell^{1}$- and $\ell^{\infty}$-operator norms:
$\|\boldsymbol{\Omega}^{(t)}\|_{1}=\max_{j}\sum_{i}\bigl|\Omega^{(t)}_{ij}\bigr|\;\leq\;2s_{\max},\qquad\|\boldsymbol{\Omega}^{(t)}\|_{\infty}=\max_{i}\sum_{j}\bigl|\Omega^{(t)}_{ij}\bigr|\;\leq\;2s_{\max}$
Finally, the standard inequality between induced norms yields
$\|\boldsymbol{\Omega}^{(t)}\|_{2}\;\leq\;\sqrt{\|\boldsymbol{\Omega}^{(t)}\|_{1}\,\|\boldsymbol{\Omega}^{(t)}\|_{\infty}}\;\leq\;\sqrt{(2s_{\max})(2s_{\max})}\;=\;2s_{\max}\;\leq\;2r$
which proves the claim. ∎
2.4 Stewart–Sun eigenvector perturbation bound
Definition 5 (Distortion factor).
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be a diagonalizable matrix with eigendecomposition
$\mathbf{A}\;=\;\mathbf{V}\boldsymbol{\Lambda}\,\mathbf{V}^{-1}$
where $\boldsymbol{\Lambda}$ is diagonal with the eigenvalues of $\mathbf{A}$ on its diagonal, and $\mathbf{V}=[\mathbf{v}_{1},\dots,\mathbf{v}_{n}]$ collects the corresponding (right) eigenvectors as columns. The distortion factor of $\mathbf{A}$ is defined by
$\kappa(\mathbf{A})\;:=\;\|\mathbf{V}\|_{2}\,\|\mathbf{V}^{-1}\|_{2}$
When the underlying matrix is clear from context, we write $\kappa$ for $\kappa(\mathbf{A})$. This quantity measures how far the eigenvector matrix $\mathbf{V}$ is from being orthogonal [^3] For directed networks, adjacency matrices need not admit an orthogonal eigenbasis. In particular, for a $\{0,1\}$ adjacency matrix $\mathbf{A}$ one has $(\mathbf{A}^{\top}\mathbf{A})_{ii}=d_{i}^{\mathrm{in}}$ and $(\mathbf{A}\mathbf{A}^{\top})_{ii}=d_{i}^{\mathrm{out}}$. Thus if $\mathbf{A}$ were (real) normal and hence orthogonally diagonalizable, $\mathbf{A}^{\top}\mathbf{A}=\mathbf{A}\mathbf{A}^{\top}$ would force $d_{i}^{\mathrm{in}}=d_{i}^{\mathrm{out}}$ for every node $i$. Therefore any node-level divergence $d_{i}^{\mathrm{in}}\neq d_{i}^{\mathrm{out}}$ rules out $\kappa(\mathbf{A})=1$ (though degree-balance alone does not guarantee $\kappa(\mathbf{A})=1$). : in particular, $\kappa(\mathbf{A})=1$ when $\mathbf{V}$ is orthogonal, i.e.,
$\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}\quad\Longleftrightarrow\quad\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle=0\ \ \text{for all }i\neq j\ \ \text{and}\ \ \|\mathbf{v}_{i}\|_{2}=1\ \ \text{for all }i$
Note that $\kappa=1$ if and only if the eigenvectors are pairwise orthogonal [^4] A somewhat rough interpretation of near pair-wise orthogonality is that the nodes that are most significant in generating one type of structure are disjoint from those that generate another, and that this is true for all structures. This would mean that, for example, the nodes most responsible for heavy tails in degree distribution have little role to play in triangles.. $\circ$
Result 1 (Stewart–Sun eigenvector perturbation bound).
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be diagonalizable with simple eigenvalues. Let $\gamma=\gamma(\mathbf{A})$ be its spectral gap (Def 3) and $\kappa=\kappa(\mathbf{A})$ its distortion factor (Def 5). $\mathbf{v}$ is the unit leading eigenvector of $\mathbf{A}$ and $\mathbf{v}^{(t)}$ the unit leading eigenvector of $\mathbf{A}^{(t)}$. Let $\theta_{\mathbf{v},\mathbf{v}^{(t)}}$ denote the acute angle between $\mathbf{v}$ and $\mathbf{v}^{(t)}$. If
$\|\boldsymbol{\Omega}^{(t)}\|_{2}\;<\;\frac{\gamma}{\kappa}$
then the Stewart and Sun (1990) perturbation bound gives
$\sin\theta_{\mathbf{v},\mathbf{v}^{(t)}}\;\leq\;\frac{\kappa\,\|\boldsymbol{\Omega}^{(t)}\|_{2}}{\gamma}$
$\circ$
The bound formalizes the idea the perturbation $\boldsymbol{\Omega}^{(t)}$ can only rotate the leading eigenvector by a small angle if the leading eigenvalue is well separated from the rest of the spectrum (large $\gamma$) and if the eigenbasis is not too ill-conditioned (small $\kappa$). In essence, $\kappa$ and (reciprocal of) $\gamma$ capture information about how petrubations to smaller eigenmodes translate to a perturbation to the leading eigenmode because of the dependencies between them. More specifically, a small spectral gap (small $\gamma$) means the top modes have nearly equal growth rates, so even perturbations that primarily affect a lower mode can readily mix with the leading mode. A large distortion factor (large $\kappa$) means the eigenvectors are far from orthogonal, so the eigenmodes are geometrically entangled: a perturbation expressed in a ‘lower’ direction can leak into the leading direction through the ill-conditioned change-of-basis. Thus $\gamma$ captures dynamic separability of modes, while $\kappa$ captures geometric separability. Naturally, either form of coupling allows disturbances to lower-ranked structures to indirectly move the leading eigenmode.
Proposition 3 (Angle of deviation between the degree vector and the eigenvector).
Let $\mathbf{A}$ be neutral (Def. 2), with unit degree vector $\mathbf{d}$ and unit leading eigenvector $\mathbf{v}$. Let $\mathbf{v}^{(t)}$ be the unit leading eigenvector of $\mathbf{A}^{(t)}=\mathbf{A}+\boldsymbol{\Omega}^{(t)}$. Assume the Stewart–Sun condition
$\|\boldsymbol{\Omega}^{(t)}\|_{2}\;<\;\frac{\gamma}{\kappa}$
where $\gamma=\gamma(\mathbf{A})$ and $\kappa=\kappa(\mathbf{A})$ are as in Result 1. Then
$\theta_{\mathbf{d},\mathbf{v}^{(t)}}\;\leq\;\theta_{\mathbf{d},\mathbf{v}}\;+\;\arcsin\Bigl(\tfrac{\kappa\,\|\boldsymbol{\Omega}^{(t)}\|_{2}}{\gamma}\Bigr)$
If, moreover, each node participates in at most $r$ swaps so that $\|\boldsymbol{\Omega}^{(t)}\|_{2}\leq 2r$ (Prop. 2), then whenever $\tfrac{2\kappa r}{\gamma}<1$
$\theta_{\mathbf{d},\mathbf{v}^{(t)}}\;\leq\;\theta_{\mathbf{d},\mathbf{v}}\;+\;\arcsin\Bigl(\tfrac{2\kappa r}{\gamma}\Bigr)$
Since $\theta_{\mathbf{d},\mathbf{v}}\approx 0$ (Prop. 1), we have
$\sin\theta_{\mathbf{d},\mathbf{v}^{(t)}}\;\lesssim\;\frac{2\kappa r}{\gamma}$
$\circ$
In other words, given our degree-preserving rewiring process, the angle of deviation between degree vector and the eigenvector is bounded from above by the maximum number of rewirings per node ($r$), a measure of the orthogonality of the eigenmodes ($\kappa$), and the reciprocal of the spectral gap ($\lambda$).
3 Bounding moments of angle of deviation with moments of network statistic
Conjecture 1 (Wandering angle under strictly $\phi$-improving rewiring).
Fix in- and out-degree sequences and let $\mathcal{S}$ be the (finite) set of directed adjacency matrices reachable from an initial $\mathbf{A}^{(0)}$ by degree-preserving single-swap rewirings. Fix a unit degree vector $\mathbf{d}$ that is invariant on $\mathcal{S}$.
Consider a degree-preserving statistic-driven rewiring process that pathwise increases a chosen statistic $\phi$ pertaining to assortativity and/or local or mesoscopic structure (e.g. triangles, $k$-cycles, communities, core–periphery). Thus, along the realized trajectory $\{\mathbf{A}^{(t)}\}_{t\geq 0}\subset\mathcal{S}$ of accepted swaps,
$\phi(\mathbf{A}^{(t+1)})\;\geq\;\phi(\mathbf{A}^{(t)})\qquad\text{for all }t\geq 0.$
For each $\mathbf{A}^{(t)}\in\mathcal{S}$, let $\mathbf{v}(\mathbf{A}^{(t)})$ be the unit leading eigenvector and let $\theta_{t}$ be the angle of deviation between $\mathbf{v}(\mathbf{A}^{(t)})$ and the fixed degree vector $\mathbf{d}$.
The conjecture is that, for a broad class of statistics $\phi$ that reward local or mesoscopic structure, the induced angle process $(\theta_{t})_{t\geq 0}$ typically does not drift with a fixed sign. That is, although $\phi(\mathbf{A}^{(t)})$ increases with $t$, the sequence $\theta_{t}$ may increase over some epochs and decrease over others. In particular, $\theta_{t}$ need not be monotone in $\phi(\mathbf{A}^{(t)})$ either pathwise or in expectation. In essence, the angle of deviation between the degree vector and the eigenvector can ‘wander’. The mechanism that generates this wandering is the tension between the two opposing forces induced by the rewiring procedure:
  1. (i)
    Pull away from degree: conditioning on $\phi(\mathbf{A}^{t+1})\geq\phi(\mathbf{A}^{t})$ biases accepted swaps toward edge rearrangements that create the local patterns rewarded by $\phi$. These patterns reallocate Perron mass across vertices without changing degrees, pushing the eigenvector $\mathbf{v}(\mathbf{A}^{(t)})$ away from the fixed degree vector $\mathbf{d}$, thereby increasing $\theta_{t}$.
  2. (ii)
    Pull toward degree: the collection of $\phi$-improving swaps is biased towards high-degree nodes. From a combinatorial point of view, there is a greater likelihood of candidate endpoints for cycles, communities, and other local structures being at or near high-degree nodes. Hence, even under uniform proposals, accepted swaps may disproportionately endow high-degree nodes with additional local structure [^5] Note that the condition in Steward–Sun that each node can participate in at most $r$ rewiring limits how much local structure can accumulate around any single vertex along the rewiring trajectory. The bounded participation condition, therefore, limits the ‘opportunity effect’ when $r$ is sufficiently small compared to the heaviness of the tails of the degree distribution.. Since $\mathbf{v}(\mathbf{A})$ frequently correlates with degree in heterogeneous graphs, this reinforcement can pull $\mathbf{v}(\mathbf{A}^{(t)})$ back toward $\mathbf{d}$ and thereby decrease $\theta_{t}$.
The interplay of (i) and (ii) can therefore produce a trajectory for $\theta_{t}$ that wanders within $0,\pi/2]$ despite the strict monotonicity of $\phi(\mathbf{A}^{(t)})$ along each step [^6] This conjecture helps explain why the empirical evidence on the matter is mixed. One line of work shows that as local or mesoscopic structure accumulates, the leading eigenvector may concentrate on a small set of vertices and drift away from degree-based rankings (Sharkey, 2019; Pastor-Satorras and Castellano, 2016). But the opposite tendency can also occur: within a fixed degree sequence, there are neutral realizations in which degree and eigenvector centrality remain close. This includes maximum-entropy/BFD graphs (Atay and Bıyıkoğlu, forthcoming). Some rewiring protocols effectively add structure in ways that track degree rather than compete with it. In short, different networks and protocols can therefore generate different outcomes. Our conjecture gives reasons for why the empirical record does not point in a single direction.
(In Def 6, we impose some structure upon the relation between the evolution of the network statistic and the angle of deviation between the degree vector and eigenvector). $\circ$
Definition 6 (Statistic-driven rewiring with mean-angle admissibility).
Let $\phi:\mathcal{S}\to\mathbb{R}$ be a network statistic and let $\mathbf{A}^{(0)}\in\mathcal{S}$ be a neutral network (Def. 2). We index the process by the number of accepted perturbations $t$. Given $\mathbf{A}^{(t)}$, let $\mathcal{N}(\mathbf{A}^{(t)})\subseteq\mathcal{S}$ denote its (degree-preserving) swap neighborhood and define the $\phi$-upper-contour set
$\mathcal{B}^{(t)}\;:=\;\bigl\{\mathbf{B}\in\mathcal{N}(\mathbf{A}^{(t)}):\ \phi(\mathbf{B})\geq\phi(\mathbf{A}^{(t)})\bigr\}$
Assume $\mathcal{B}_{t}\neq\emptyset$ along the horizon of interest [^7] We work in regimes with only a limited amount of assortativity tuning and a modest accumulation of local structures, far from saturation, so the $\phi$-upper contour remains nontrivial along the realized trajectory.. Let $\widetilde{\mathcal{B}}_{t}\subseteq\mathcal{B}_{t}$ be a subset satisfying the mean-angle admissibility condition
$\mathbb{E}\bigl[\theta(\mathbf{A}^{(t+1)})\mid\mathbf{A}^{(t)}\bigr]\ \geq\ \theta(\mathbf{A}^{(t)}),\qquad\text{with }\mathbf{A}^{(t+1)}\in\widetilde{\mathcal{B}}^{(t)}t$
In general, there may be many such admissible subsets. We take $\widetilde{\mathcal{B}}^{t}$ to be the largest (by inclusion) subset of $\mathcal{B}^{t}$ for which the above inequality holds. [^8] This is weaker than requiring $\theta(\mathbf{A}^{(t+1)})\geq\theta(\mathbf{A}^{(t)})$ for every accepted move. Imposing only nonnegative drift in expectation typically makes the admissible set much less likely to be empty when $\mathcal{B}_{t}$ is large. Finally, choose $\mathbf{A}^{(t+1)}\in\widetilde{\mathcal{B}}_{t}$ according to a sampling rule on $\widetilde{\mathcal{B}}_{t}$. Define $\phi_{t}:=\phi(\mathbf{A}^{(t)})$. By construction, $\phi_{t+1}\geq\phi_{t}$ along every realized trajectory and $\theta_{t}$ is a submartingale. $\circ$
Note that the angle-biased selection rule above is a deliberately conservative modeling choice: we explicitly allow the degree–eigenvector angle to deviate as the rewiring procedure builds up assortativity or other local structures. Consequently, the bounds we derive in later sections on how much $\theta(\mathbf{A})$ can change as a function of assortativity (or other structure-rewarding statistics) are upper bounds: they quantify the maximal misalignment compatible with the imposed statistic trajectory and the rewiring budget, not a typical or inevitable misalignment along every realization.
Proposition 4 (Existence and monotonicity of moments for a bounded network statistic).
Let $\{\mathbf{A}^{(t)}\}_{t\geq 0}$ be the network statistic driven degree-preserving rewiring process from Def 6. Assume that $\phi$ is bounded on the state space reachable by the rewiring, i.e., there exist constants $\underline{\phi}<\overline{\phi}$ such that
$\underline{\phi}\leq\phi(\mathbf{A})\leq\overline{\phi}\qquad\text{for all reachable }\mathbf{A}$
Then for every $t\geq 0$ and every $p\geq 1$ the moment $\mathbb{E}[|\phi_{t}|^{p}]$ is well-defined and finite. In particular, $\mathbb{E}[\phi_{t}]$ and $\mathbb{E}[\phi_{t}^{2}]$ exist for all $t$.
If, in addition, $\phi$ is strictly positive on the reachable state space (e.g., $0<\phi(\mathbf{A})\leq\overline{\phi}$), then for every $p\geq 1$ the $p$th moment is increasing in $t$:
$\mathbb{E}[\phi_{t+1}^{p}]\;\geq\;\mathbb{E}[\phi_{t}^{p}]\qquad\text{for all }t\geq 0$
Indeed, $t$ counts accepted swaps, so the acceptance rule enforces the pathwise increase $\phi_{t+1}\geq\phi_{t}$ almost surely. Under positivity, the map $x\mapsto x^{p}$ is increasing on $\mathbb{R}_{>0}$, hence $\phi_{t+1}^{p}\geq\phi_{t}^{p}$ almost surely, and taking expectations yields the claim. In particular,
$\mathbb{E}[\phi_{t+1}]\;\geq\;\mathbb{E}[\phi_{t}],\qquad\mathbb{E}[\phi_{t+1}^{2}]\;\geq\;\mathbb{E}[\phi_{t}^{2}],\qquad\text{for all }t\geq 0$
Note that even under positivity and pathwise increase, the variance $\mathbb{V}(\phi_{t})$ need not be monotone in $t$: while both $\mathbb{E}[\phi_{t}]$ and $\mathbb{E}[\phi_{t}^{2}]$ increase, the difference $\mathbb{V}(\phi_{t})=\mathbb{E}[\phi_{t}^{2}]-\bigl(\mathbb{E}[\phi_{t}]\bigr)^{2}$ can increase or decrease depending on their relative rates of growth. In fact, in general, $\mathbb{V}(\phi_{t})$ will tend to increase in the early stages of rewiring as more and diverse networks become ‘reachable’ within one step. After a sufficiently large number of rewirings, $\mathbb{V}(\phi_{t})$ will tend to decrease as the statistic reaches the neighborhood of its upper bound, and therefore networks reachable within one step tend not to be able to offer much of an improvement. $\circ$
Proposition 5 (Moment transfer via local linearization).
Let $\{\mathbf{A}^{(t)}\}_{t\geq 0}$ be an $\mathcal{S}$-valued rewiring process with
$\mathbf{A}^{(t)}=\mathbf{A}^{(0)}+\boldsymbol{\Omega}^{(t)},\qquad\boldsymbol{\Omega}^{(t)}=\sum_{s=0}^{t-1}\boldsymbol{\Delta}^{(s)}$
where each swap induces a low-rank perturbation $\boldsymbol{\Delta}^{(s)}$ with uniformly bounded size (so the steps are marginal in operator norm). Let
$\theta_{t}:=\theta(\mathbf{A}^{(t)})\in[0,\pi/2],\qquad\phi_{t}:=\phi(\mathbf{A}^{(t)})$
for a network statistic $\phi:\mathcal{S}\to\mathbb{R}$.
Assume that over the horizon of interest the trajectory remains in a subset $\mathcal{S}_{0}\subseteq\mathcal{S}$ on which the leading eigenvalue is well separated and the eigenbasis is not too ill-conditioned. In particular, on any reachable set $\mathcal{S}_{0}$ supporting the affine comparison below one may take uniform spectral parameters [^9] The reachable region $\mathcal{S}_{0}$ (hence $\kappa_{*},\gamma_{*}$) may depend on the statistic $\phi$ and the acceptance rule, since different driving statistics can steer the trajectory through regions with different spectral conditioning.
$\kappa_{*}:=\sup_{\mathbf{A}\in\mathcal{S}_{0}}\kappa(\mathbf{A}),\qquad\gamma_{*}:=\inf_{\mathbf{A}\in\mathcal{S}_{0}}\gamma(\mathbf{A})\;>\;0$
On such a region, eigenvector perturbation theory (Stewart–Sun) implies that $\mathbf{A}\mapsto\theta(\mathbf{A})$ is locally Lipschitz in the operator norm: there exists $L_{\theta}>0$ such that for all $\mathbf{A},\mathbf{B}\in\mathcal{S}_{0}$,
$|\theta(\mathbf{B})-\theta(\mathbf{A})|\;\leq\;L_{\theta}\,\|\mathbf{B}-\mathbf{A}\|_{2},\qquad\text{with }L_{\theta}\asymp\kappa_{*}/\gamma_{*}$
Assume further that, along $\mathcal{S}_{0}$, the statistic $\phi$ tracks the same marginal edits in the sense that there exist constants $a^{\pm}\geq 0$, $b^{\pm}\in\mathbb{R}$, and a finite-size slack $\varepsilon_{n}\geq 0$ such that the following affine bounds hold:
$a^{-}\,\phi(\mathbf{A})+b^{-}-\varepsilon_{n}\ \leq\ \theta(\mathbf{A})\ \leq\ a^{+}\,\phi(\mathbf{A})+b^{+}+\varepsilon_{n},\qquad\forall\,\mathbf{A}\in\mathcal{S}_{0}$
Here the effective constants $(a^{\pm},b^{\pm},\varepsilon_{n})$ may depend on the uniform spectral controls $(\kappa_{*},\gamma_{*})$ (through the stability of $\theta(\cdot)$ under marginal edits) and on statistic-specific parameters that govern how $\phi(\cdot)$ responds to an accepted swap (e.g. partition granularity, $d_{\max}$-type caps, motif size, or acceptance-rule constraints).
If $\phi$ is bounded on the reachable state space (so that $\mathbb{E}[\phi_{t}]$ and $\mathbb{E}[\phi_{t}^{2}]$ exist), then for every $t$
$a^{-}\,\mathbb{E}[\phi_{t}]+b^{-}-\varepsilon_{n}\ \leq\ \mathbb{E}[\theta_{t}]\ \leq\ a^{+}\,\mathbb{E}[\phi_{t}]+b^{+}+\varepsilon_{n}$
Moreover, with $c^{+}:=|b^{+}|+\varepsilon_{n}$
$\mathbb{E}[\theta_{t}^{2}]\ \leq\ (a^{+})^{2}\,\mathbb{E}[\phi_{t}^{2}]\;+\;2a^{+}c^{+}\,\mathbb{E}[|\phi_{t}|]\;+\;(c^{+})^{2}$
$\circ$
Corollary 1 (Multiplicative envelope for signed statistics).
Fix a reachable region $S_{0}$ and a statistic $\phi:\mathcal{S}\to\mathbb{R}$. Suppose the affine upper comparison from Proposition 5 holds on $S_{0}$:
$\theta(A)\;\leq\;a_{\phi}^{+}\,\phi(A)+b_{\phi}^{+}+\varepsilon_{n},\qquad\forall A\in S_{0}$
Then, for all $A\in S_{0}$,
$\theta(A)\;\leq\;a_{\phi}^{+}\,|\phi(A)|+\underbrace{\bigl(|b_{\phi}^{+}|+\varepsilon_{n}\bigr)}_{=:c_{\phi}}$
Let $M_{\phi}:=\max\{a_{\phi}^{+},c_{\phi}\}$. Since $a_{\phi}^{+}|\phi(A)|+c_{\phi}\leq M_{\phi}|\phi(A)|+M_{\phi}$, we obtain the purely algebraic multiplicative envelope
$\theta(A)\;\leq\;M_{\phi}\bigl(1+|\phi(A)|\bigr),\qquad\forall A\in S_{0}$
and hence along the trajectory
$\mathbb{E}[\theta_{t}]\;\leq\;M_{\phi}\Bigl(1+\mathbb{E}\bigl[|\phi_{t}|\bigr]\Bigr)$
(Chaining linear bounds to obtain a product envelope for $M_{\phi}$.) Assume that on $S_{0}$ the following linear controls hold:
(i) Linear spectral sensitivity (Stewart–Sun). There is a constant $L_{\theta}\lesssim\kappa_{*}/\gamma_{*}$ such that for all $A\in S_{0}$,
$\theta(A)\;\leq\;\theta(A^{(0)})+L_{\theta}\,\|A-A^{(0)}\|_{2}$
(ii) Participation-controlled perturbation size (no explicit $n$). Under $r_{\phi}$-bounded participation, the cumulative perturbation $\Omega^{(t)}=\sum_{s=0}^{t-1}\Delta^{(s)}$ satisfies
$\|\Omega^{(t)}\|_{2}\;\leq\;\sqrt{\|\Omega^{(t)}\|_{1}\,\|\Omega^{(t)}\|_{\infty}}\;\leq\;c_{\Omega}\,r_{\phi}$
where $\|\cdot\|_{\infty}$ and $\|\cdot\|_{1}$ are the maximum absolute row- and column-sum norms. Here $c_{\Omega}>0$ is universal (e.g. $c_{\Omega}=2$ under the two-edge swap model), because $r_{\phi}$-bounded participation bounds the total signed mass that can accumulate in any fixed row or column of $\Omega^{(t)}$.
(iii) Linear leverage of the driving statistic. There is a statistic-specific scale $\Lambda_{\phi}$ such that each accepted swap changes the statistic by at most $c_{\phi}^{\mathrm{lev}}\Lambda_{\phi}$:
$|\phi_{t+1}-\phi_{t}|\;\leq\;c_{\phi}^{\mathrm{lev}}\Lambda_{\phi}$
with $c_{\phi}^{\mathrm{lev}}$ universal. [^10] As before, this is a bounded-differences property: a single swap alters only $O(1)$ edge incidences, so $|\Delta\phi|$ is controlled by degree caps, partition granularity, motif length, and the normalization of $\phi$. In normalized statistics, $\Lambda_{\phi}$ typically already carries factors such as $1/m$, $1/|C_{k}|^{2}$, etc., which prevents spurious growth with $n$.
Chaining (i) and (ii) gives the basic linear envelope
$\theta(A^{(t)})\;\leq\;\theta(A^{(0)})+\Bigl(\frac{\kappa_{*}}{\gamma_{*}}\Bigr)\,(c_{\Omega}r_{\phi})$
Combining this with (iii) (to express the same perturbation budget in the natural leverage scale of the driving statistic) yields the product-form envelope [^11] Whenever the ingredients are linear in their respective controls, chaining them automatically produces a multiplicative envelope: if $u\leq c_{1}x$, $x\leq c_{2}y$, and $y\leq c_{3}z$ (with nonnegative quantities), then substitution gives $u\leq(c_{1}c_{2}c_{3})\,z$. In the present setting one may read the chain schematically as $\theta-\theta_{0}\;\leq\;L_{\theta}\,\|\Omega\|_{2},\qquad L_{\theta}\;\lesssim\;\frac{\kappa_{*}}{\gamma_{*}},\qquad\|\Omega\|_{2}\;\leq\;c_{\Omega}\,r_{\phi},$ together with the statistic-side leverage scale $\Lambda_{\phi}$ that governs the affine comparison constants (via degree caps, partition granularity, motif length, and normalization). Chaining these linear bounds yields a product of the contributing factors, while additive intercepts and finite-size slack are absorbed into the prefactor $C_{\phi}$.
$M_{\phi}\;\lesssim\;C_{\phi}\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\phi}\,\Lambda_{\phi}$
where $C_{\phi}$ absorbs baseline terms ($\theta(A^{(0)})$, $|\phi_{0}|$), universal constants, the normalization conventions of $\phi$, and any residual slack from the affine comparison [^12] The product form is a bookkeeping consequence of chaining linear inequalities: each step contributes a linear factor (spectral sensitivity, perturbation size, statistic leverage), and their product dominates the final upper envelope once constants are taken large enough to absorb additive slack..
4 Assortativity
Definition 7 (Newman’s assortativity coefficient).
Let $\mathbf{A}$ be the adjacency matrix of a (possibly directed) graph with edge set $\mathcal{E}(\mathbf{A})$ and $m:=|\mathcal{E}(\mathbf{A})|$. For each vertex $i$, let $d_{i}^{\mathrm{in}}$ and $d_{i}^{\mathrm{out}}$ denote its in- and out-degree. Fix $(p,q)\in\{\mathrm{in},\mathrm{out}\}^{2}$. For each directed edge $(i,j)\in\mathcal{E}(\mathbf{A})$ define the tail and head degrees
$x_{ij}:=d_{i}^{p},\qquad y_{ij}:=d_{j}^{q}.$
Let $\mu_{T}:=\frac{1}{m}\sum_{(i,j)\in\mathcal{E}(\mathbf{A})}x_{ij}$ and $\mu_{H}:=\frac{1}{m}\sum_{(i,j)\in\mathcal{E}(\mathbf{A})}y_{ij}$, and let
$\sigma_{T}^{2}:=\frac{1}{m}\sum_{(i,j)\in\mathcal{E}(\mathbf{A})}(x_{ij}-\mu_{T})^{2},\qquad\sigma_{H}^{2}:=\frac{1}{m}\sum_{(i,j)\in\mathcal{E}(\mathbf{A})}(y_{ij}-\mu_{H})^{2}$
Newman’s assortativity coefficient for the choice $(p,q)$ is the Pearson correlation of $(x_{ij},y_{ij})$ over edges
$\phi_{p,q}(\mathbf{A})\;:=\;\frac{\frac{1}{m}\sum_{(i,j)\in\mathcal{E}(\mathbf{A})}(x_{ij}-\mu_{T})(y_{ij}-\mu_{H})}{\sigma_{T}\sigma_{H}}$
$\circ$
Proposition 6 (Change in Newman’s assortativity under degree-preserving rewiring).
Fix $(p,q)\in\{\mathrm{in},\mathrm{out}\}^{2}$ and let $\phi_{p,q}(\cdot)$ denote Newman’s assortativity coefficient from Def 7. Let $\mathbf{A}^{(0)}$ be a directed adjacency matrix with fixed in- and out-degree sequences, and let $\mathbf{A}^{(t)}$ be any matrix reachable from $\mathbf{A}^{(0)}$ by degree-preserving edge swaps (so $\mathbf{A}^{(t)}\in\mathcal{S}$ in the notation of Conjecture 1). Then the edge-averaged means and variances entering $\phi_{p,q}$ are invariants of the degree class:
$\mu_{T}(\mathbf{A}^{(t)})=\mu_{T}(\mathbf{A}^{(0)}),\qquad\mu_{H}(\mathbf{A}^{(t)})=\mu_{H}(\mathbf{A}^{(0)})$
$\sigma_{T}(\mathbf{A}^{(t)})=\sigma_{T}(\mathbf{A}^{(0)}),\qquad\sigma_{H}(\mathbf{A}^{(t)})=\sigma_{H}(\mathbf{A}^{(0)})$
and hence the normalization factor
$\nu_{p,q}(\mathbf{A}^{(t)})=\frac{1}{\sigma_{T}(\mathbf{A}^{(t)})\,\sigma_{H}(\mathbf{A}^{(t)})}$
is constant along the rewiring trajectory. Denote this constant by $\nu_{p,q}$. Consequently, the change in assortativity along any degree-preserving rewiring is entirely driven by the change in the edge-wise degree-product sum
$\mathcal{S}_{p,q}(\mathbf{A})\;:=\;\sum_{(i,j)\in\mathcal{E}(\mathbf{A})}d_{i}^{p}\,d_{j}^{q}$
namely,
$\phi_{p,q}(\mathbf{A}^{(t)})-\phi_{p,q}(\mathbf{A}^{(0)})\;=\;\nu_{p,q}\left[\frac{1}{m}\,\mathcal{S}_{p,q}(\mathbf{A}^{(t)})\;-\;\frac{1}{m}\,\mathcal{S}_{p,q}(\mathbf{A}^{(0)})\right]$
with $m:=|\mathcal{E}(\mathbf{A}^{(0)})|$. In particular, degree-preserving rewiring leaves unchanged the multiset of tail degrees $\{d_{i}^{p}:(i,j)\in\mathcal{E}(\mathbf{A})\}$ and head degrees $\{d_{j}^{q}:(i,j)\in\mathcal{E}(\mathbf{A})\}$. It only changes how tails and heads are paired across edges. The constant $\nu_{p,q}$ therefore acts as a fixed conversion factor from changes in the average edge-wise product $\mathcal{S}_{p,q}(\mathbf{A})/m$ to changes in Newman’s assortativity coefficient. $\circ$
Note that the conversion factor depends on degree heterogeneity. Recall $\nu_{p,q}=1/(\sigma_{T}\sigma_{H})$, where $\sigma_{T}^{2}$ and $\sigma_{H}^{2}$ are the edge-averaged variances of tail and head degrees used in Def 7. Thus $\nu_{p,q}$ is large when either $\sigma_{T}$ or $\sigma_{H}$ is small, and it is small only when both $\sigma_{T}$ and $\sigma_{H}$ are large. Intuitively, if one side is nearly homogeneous (say $\sigma_{H}\approx 0$) while the other is heterogeneous, then small changes in how the heterogeneous degrees are matched can produce comparatively large changes in the Pearson correlation, so assortativity moves quickly. By contrast, when both ends of edges are highly heterogeneous (large $\sigma_{T}$ and large $\sigma_{H}$), the background variability is large and a fixed change in the mean product $\mathcal{S}_{p,q}(\mathbf{A})/m$ translates into a smaller change in $\phi_{p,q}$.
Assortativity inducing rewiring
Fix $(p,q)\in\{\mathrm{in},\mathrm{out}\}^{2}$ and let $\phi:=\phi_{p,q}$ be Newman’s assortativity coefficient (Def. 7). Consider the statistic-driven rewiring rule of Def. 6 with this choice of $\phi$. Along the induced trajectory $\{\mathbf{A}^{(t)}\}_{t\geq 0}$, a candidate move from $\mathbf{A}^{(t)}$ to $\mathbf{A}^{(t+1)}=\mathbf{A}^{(t)}+\boldsymbol{\Delta}^{(t)}$ is a degree-preserving edge swap (Def. 6), removing $(a,b)$ and $(c,d)$ and adding $(a,d)$ and $(c,b)$ (with the usual feasibility conditions). Define the degree-product sum
$\mathcal{S}_{p,q}(\mathbf{A})\;:=\;\sum_{(i,j)\in\mathcal{E}(\mathbf{A})}d_{i}^{p}\,d_{j}^{q}$
For the above swap one has
$\mathcal{S}_{p,q}(\mathbf{A}^{(t+1)})-\mathcal{S}_{p,q}(\mathbf{A}^{(t)})\;=\;d_{a}^{p}d_{d}^{q}+d_{c}^{p}d_{b}^{q}-d_{a}^{p}d_{b}^{q}-d_{c}^{p}d_{d}^{q}$
Moreover, along any degree-preserving trajectory the normalization in $\phi_{p,q}$ is invariant (Prop. 6), so
$\phi_{p,q}(\mathbf{A}^{(t+1)})-\phi_{p,q}(\mathbf{A}^{(t)})\quad\text{has the same sign as}\quad\mathcal{S}_{p,q}(\mathbf{A}^{(t+1)})-\mathcal{S}_{p,q}(\mathbf{A}^{(t)}).$
Accordingly, a swap at time $t$ is called positive-assortativity inducing if $\mathcal{S}_{p,q}(\mathbf{A}^{(t+1)})>\mathcal{S}_{p,q}(\mathbf{A}^{(t)})$, and negative-assortativity inducing if the inequality is reversed. Note that since our assortativity moves are degree-preserving, Proposition 3 applies verbatim: under $r$-bounded participation and a neutral baseline, $\sin\theta_{\mathbf{d},\mathbf{v}^{(t)}}\leq 2\kappa r/\gamma$ (for $2\kappa r/\gamma<1$), so degree remains a good proxy as long as $r\ll\gamma/\kappa$ [^13] Instead of bounding each node’s participation by $r$, one may bound rewiring at the level of groups: partition vertices into groups of size $m$ and require $\sum_{u\in G_{\ell}}s(u)\leq R$ for each group. Since then $s_{\max}\leq R$, we have $\|\Omega^{(t)}\|_{2}\leq 2R$ and hence $\sin\theta_{\mathbf{d},\mathbf{v}^{(t)}}\leq 2\kappa R/\gamma$ (when $2\kappa R/\gamma<1$). Writing $R=rm$ shows the bound scales with group size: larger $m$ weakens the worst-case guarantee because a fixed group budget can be concentrated on fewer nodes..
Remark 1 (Moment bound for the angle via assortativity).
Let $\phi_{p,q}(\mathbf{A})$ be the Newman assortativity statistic from Def 7, $(p,q)\in\{\mathrm{in},\mathrm{out}\}^{2}$, and along the degree-preserving assortativity rewiring trajectory set
$\phi_{t}:=\phi_{p,q}(\mathbf{A}^{(t)}),\qquad\theta_{t}:=\theta(\mathbf{A}^{(t)})$
The rule is $\phi_{p,q}$-upper-contour (accepted swaps satisfy $\phi_{t+1}\geq\phi_{t}$). If the affine upper comparison assumed in Proposition 5 holds on the reachable set for $\phi=\phi_{p,q}$, then Corollary 1 implies the one-moment control
$\mathbb{E}[\theta_{t}]\;\leq\;M_{p,q}\Bigl(1+\mathbb{E}\bigl[|\phi_{t}|\bigr]\Bigr)$
for a uniform constant $M_{p,q}$.
(Assortativity-specific scaling of $M_{p,q}$.) Recall that $\phi_{p,q}$ is a normalized covariance-type statistic between the $p$-degree at the tail and the $q$-degree at the head of a directed edge, with normalization
$\nu_{p,q}=\frac{1}{\sigma_{T}\sigma_{H}}$
Under degree-preserving rewiring the degree sequence (hence $\sigma_{T},\sigma_{H}$ and the maxima $d_{\max}^{p}:=\max_{i}d_{i}^{p}$, $d_{\max}^{q}:=\max_{i}d_{i}^{q}$) remain fixed. A single swap replaces only $O(1)$ directed edges, so the assortativity numerator changes by a sum of $O(1)$ degree-products, each bounded by $d_{\max}^{p}d_{\max}^{q}$; after normalization, the one-step variation scale is therefore
$\Lambda_{p,q}\asymp\nu_{p,q}\,d_{\max}^{p}d_{\max}^{q}$
up to universal constants (and the precise convention used in Def 7). Combining this leverage estimate with the perturbation-theoretic factor $\kappa_{*}/\gamma_{*}$ and the rewiring budget/participation parameter $r$ (as it enters the operator-norm control of the cumulative perturbation) gives the multiplicative envelope
$M_{p,q}\;\lesssim\;C_{p,q}\,\frac{\kappa_{*}}{\gamma_{*}}\,r\,\Lambda_{p,q}$
where $C_{p,q}$ is chosen large enough to absorb normalization conventions and any residual slack terms from the affine comparison.
If the in/out maxima are comparable and the common degree tail is power-law with exponent $\alpha>1$ (so $d_{\max}\asymp n^{1/\alpha}$ up to slowly varying factors), then $\Lambda_{p,q}\asymp\nu_{p,q}d_{\max}^{2}$ and hence
$M_{p,q}\;\lesssim\;\widetilde{C}_{p,q}\,\nu_{p,q}\,\frac{\kappa_{*}}{\gamma_{*}}\,r\,n^{2/\alpha}$
so heavier tails (smaller $\alpha$) enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ through increased extremal-degree leverage. $\circ$
5 Local structures
We now turn to local and mesoscopic structures that can pull the leading eigenvector away from the degree proxy. We focus on three canonical patterns: communities (modularity), core-periphery organization, and cycles. For each pattern we specify a degree-preserving rewiring rule that amplifies the structure, express the resulting change as a cumulative perturbation of the adjacency matrix, and then invoke the Stewart-Sun perturbation bound (Result 1), together with our operator-norm control of the perturbation, to bound the induced degree-eigenvector deviation.
5.1 Communities
Definition 8 (Community-contrast statistic).
Let $\mathbf{A}\in\{0,1\}^{n\times n}$ be a directed simple graph with $m:=|\mathcal{E}(\mathbf{A})|$ edges, and fix a partition $\mathcal{C}=\{C_{1},\dots,C_{K}\}$. Define the block edge fractions
$e_{k\ell}\;:=\;\frac{1}{m}\sum_{i\in C_{k}}\sum_{j\in C_{\ell}}a_{ij},\qquad 1\leq k,\ell\leq K$
and the corresponding block marginals
$e_{k\cdot}:=\sum_{\ell=1}^{K}e_{k\ell},\qquad e_{\cdot k}:=\sum_{\ell=1}^{K}e_{\ell k}$
The community-contrast of $\mathbf{A}$ with respect to $\mathcal{C}$ is
$\phi_{\mathrm{com}}(\mathbf{A};\mathcal{C})\;:=\;\sum_{k=1}^{K}\bigl(e_{kk}-e_{k\cdot}\,e_{\cdot k}\bigr)$
This statistic is positive when within-community edge mass exceeds the baseline predicted by the block marginals, and is small (or negative) when edges are predominantly between communities. $\circ$
Community strengthening rewiring
Let $\mathbf{A}^{(0)}$ be a neutral baseline (Def 2) and fix a partition of the vertex set into two groups $\mathcal{Q}_{1}\cup\mathcal{Q}_{2}=\{1,\dots,n\}$, chosen independently of degrees. Impose the angle deviation constraint in Def 6. A community-forming swap selects two cross-community edges $(a,b)\in\mathcal{E}_{12}$ and $(c,d)\in\mathcal{E}_{21}$ and performs the standard degree-preserving swap
[…]
thereby converting two cross edges into one within-$\mathcal{Q}_{1}$ edge and one within-$\mathcal{Q}_{2}$ edge while preserving all in- and out-degrees.
Writing $\boldsymbol{\Delta}^{(i)}_{\mathrm{com}}$ for the perturbation matrix of the $i$th such swap and $\boldsymbol{\Omega}_{\mathrm{com}}^{(t)}:=\sum_{i=1}^{t}\boldsymbol{\Delta}^{(i)}_{\mathrm{com}}$, the rewired adjacency matrix is $\mathbf{A}^{(t)}=\mathbf{A}^{(0)}+\boldsymbol{\Omega}_{\mathrm{com}}^{(t)}$. If each vertex participates in at most $r_{\mathrm{com}}$ community-forming swaps, then Proposition 2 gives
$\bigl\|\boldsymbol{\Omega}_{\mathrm{com}}^{(t)}\bigr\|_{2}\;\leq\;2r_{\mathrm{com}}$
Hence, whenever $2\kappa r_{\mathrm{com}}/\gamma<1$, Proposition 3 along with the neutrality assumption yields the eigenvector-rotation bound
$\sin\theta_{\mathbf{d},\mathbf{v}^{(t)}}\;\leq\;\frac{2\kappa r_{\mathrm{com}}}{\gamma}$
Extension to many communities
The same construction applies to any partition $\mathcal{Q}_{1},\dots,\mathcal{Q}_{m}$ with $m\geq 2$: pick any pair of groups $(p,q)$, choose two edges in opposite directions between $\mathcal{Q}_{p}$ and $\mathcal{Q}_{q}$, and swap them into within-$\mathcal{Q}_{p}$ and within-$\mathcal{Q}_{q}$ edges. All bounds above remain unchanged, since they depend only on degree preservation and the participation budget $r_{\mathrm{com}}$ through Proposition 2 and Proposition 3.
Remark 2 (Moment bound for the angle via the community statistic).
Let $\phi_{\mathrm{com}}(\mathbf{A};\mathcal{C})$ be the community-contrast statistic from Def 8, computed for a fixed partition $\mathcal{C}=\{C_{1},\dots,C_{K}\}$. Along the degree-preserving community-forming rewiring trajectory $\{\mathbf{A}^{(t)}\}_{t\geq 0}$ define
$\phi_{t}:=\phi_{\mathrm{com}}(\mathbf{A}^{(t)};\mathcal{C}),\qquad\theta_{t}:=\theta(\mathbf{A}^{(t)}).$
The acceptance rule is $\phi_{\mathrm{com}}$-upper-contour (accepted swaps satisfy $\phi_{t+1}\geq\phi_{t}$). If the affine upper comparison assumed in Proposition 5 holds on the reachable set for $\phi=\phi_{\mathrm{com}}(\cdot;\mathcal{C})$, then Corollary 1 implies the one-moment control
$\mathbb{E}[\theta_{t}]\;\leq\;M_{\mathrm{com}}\Bigl(1+\mathbb{E}\bigl[|\phi_{t}|\bigr]\Bigr),$
for a uniform constant $M_{\mathrm{com}}$.
(Community-specific scaling of $M_{\mathrm{com}}$.) The partition enters only through its granularity $K$ and block sizes $\{|C_{k}|\}_{k=1}^{K}$. Under degree-preserving swaps the degree sequence is fixed, so the relevant extremal controls are $d_{\max}^{\mathrm{out}}$ and $d_{\max}^{\mathrm{in}}$. A single community-forming swap replaces $O(1)$ edges and can change the within-block edge count only by $O(1)$; in weighted form, the largest mass a swap can redirect into within-community blocks is controlled by the extremal degree product $d_{\max}^{\mathrm{out}}d_{\max}^{\mathrm{in}}$ (up to the normalization used in Def 8). Thus the natural degree-leverage scale is
$\Lambda_{\mathrm{com}}\;\asymp\;C_{\mathrm{com}}(\mathcal{C})\,d_{\max}^{\mathrm{out}}d_{\max}^{\mathrm{in}},$
where $C_{\mathrm{com}}(\mathcal{C})$ collects the partition-dependent normalization (e.g. factors depending on $K$ and $\{|C_{k}|\}$). Combining this leverage estimate with the perturbation factor $\kappa_{*}/\gamma_{*}$ and the community participation/budget parameter $r_{\mathrm{com}}$ gives the multiplicative envelope
$M_{\mathrm{com}}\;\lesssim\;\widetilde{C}_{\mathrm{com}}(\mathcal{C})\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\mathrm{com}}\,\Lambda_{\mathrm{com}},$
with $\widetilde{C}_{\mathrm{com}}(\mathcal{C})$ chosen large enough to absorb normalization conventions and any residual slack terms from the affine comparison.
If the in/out maxima are comparable and the common degree tail is power-law with exponent $\alpha>1$ (so $d_{\max}^{\mathrm{out}}\asymp d_{\max}^{\mathrm{in}}\asymp d_{\max}\asymp n^{1/\alpha}$ up to slowly varying factors), then $\Lambda_{\mathrm{com}}\asymp C_{\mathrm{com}}(\mathcal{C})\,d_{\max}^{2}$ and hence
$M_{\mathrm{com}}\;\lesssim\;\widehat{C}_{\mathrm{com}}(\mathcal{C})\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\mathrm{com}}\,n^{2/\alpha},$
so heavier tails (smaller $\alpha$) enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ through increased extremal-degree leverage under degree-preserving swaps. $\circ$
5.2 Core-periphery
Definition 9 (Core-periphery contrast statistic).
Let $\mathbf{A}\in\{0,1\}^{n\times n}$ be a directed simple graph with $m:=|\mathcal{E}(\mathbf{A})|$ edges, and fix a core-periphery partition $\{1,\dots,n\}=\mathcal{H}\sqcup\mathcal{L}$. For $X,Y\in\{\mathcal{H},\mathcal{L}\}$ define the block edge fractions
$e_{XY}\;:=\;\frac{1}{m}\sum_{i\in X}\sum_{j\in Y}a_{ij}$
The core-periphery contrast of $\mathbf{A}$ (relative to $\mathcal{H},\mathcal{L}$) is
$\phi_{\mathrm{cp}}(\mathbf{A};\mathcal{H},\mathcal{L})\;:=\;1-2e_{\mathcal{L}\mathcal{L}}\;=\;\bigl(e_{\mathcal{H}\mathcal{H}}+e_{\mathcal{H}\mathcal{L}}+e_{\mathcal{L}\mathcal{H}}\bigr)-e_{\mathcal{L}\mathcal{L}}$
Thus $\phi_{\mathrm{cp}}$ is large when within-periphery density is small, and it decreases as $\mathcal{L}\to\mathcal{L}$ edges accumulate. $\circ$
Degree-based partition and a swap
We implement a degree-preserving rewiring that strengthens a core-periphery pattern: a ‘core’ $\mathcal{H}$ that is dense internally and well connected to a ‘periphery’ $\mathcal{L}$, with relatively few $\mathcal{L}\to\mathcal{L}$ links. Starting from a neutral baseline $\mathbf{A}$ (Def 2), order vertices by (out-)degree and split them into a high-degree set $\mathcal{H}$ (core) and a low-degree set $\mathcal{L}$ (periphery) [^14] This degree-biased choice is essential because the hub by definition has more connections than the periphery. The hub must therefore have high-degree nodes as our rewiring process preserves degree.. A single core-periphery-forming step is the degree-preserving swap that replaces one periphery-periphery edge with a core-periphery edge: pick
$(a\to b)\ \text{with }a,b\in\mathcal{L},\qquad(c\to d)\ \text{with }c\in\mathcal{H}$
delete these two edges, and add
$(a\to d),\qquad(c\to b)$
whenever this creates no duplicate edges. This removes an $\mathcal{L}\to\mathcal{L}$ link and forces an $\mathcal{H}\to\mathcal{L}$ link, thereby reducing within-periphery density and strengthening core-to-periphery connectivity. Note that $d$ can belong to either $\mathcal{L}$ or $\mathcal{H}$, which is precisely what ensures that the density of connections within the hub does not decline as the rewiring process unfolds.
Perturbation and eigenvector control
Let $\boldsymbol{\Delta}^{(t)}_{cp}$ be the perturbation matrix of the $t^{\mathrm{th}}$ core-periphery swap, and set
$\boldsymbol{\Omega}^{(t)}_{\mathrm{cp}}:=\sum_{s=1}^{t}\boldsymbol{\Delta}^{(s)}_{\mathrm{cp}},\qquad\mathbf{A}^{(t)}=\mathbf{A}^{(0)}+\boldsymbol{\Omega}^{(t)}_{\mathrm{cp}}$
Under an $r_{\mathrm{cp}}$-bounded participation budget, Proposition 2 gives $\|\boldsymbol{\Omega}^{(t)}_{\mathrm{cp}}\|_{2}\leq 2r_{\mathrm{cp}}$. Hence, whenever $2\kappa r_{\mathrm{cp}}/\gamma<1$, Proposition 3 yields
$\sin\bigl(\theta_{\mathbf{d},\mathbf{v}^{(t)}}\bigr)\ \leq\ \frac{2\kappa r_{\mathrm{cp}}}{\gamma}$
Since $\mathbf{A}^{(t)}$ shares the same in- and out-degree sequences as the neutral baseline, Proposition 1 then transfers this control to the degree–eigenvector misalignment: for modest $r_{\mathrm{cp}}$ (relative to $\gamma/\kappa$), degree remains a reliable proxy even as the rewiring amplifies a core-periphery pattern.
Fractal core-periphery generalization
The same degree-preserving swap can be iterated on nested partitions to create a hierarchical core-periphery pattern. Start with the degree-based split $\{1,\dots,n\}=\mathcal{H}^{(0)}\cup\mathcal{L}^{(0)}$. Apply the core-periphery-forming rewiring to strengthen connectivity within $\mathcal{H}^{(0)}$ and from $\mathcal{H}^{(0)}$ to $\mathcal{L}^{(0)}$, producing $\mathbf{A}^{(t)}=\mathbf{A}^{(0)}+\boldsymbol{\Omega}^{(0)}$ after the level-0 swaps.
To generate additional levels, recursively refine the periphery: at level $\ell\geq 1$, partition each periphery block from level $\ell-1$ into sub-blocks, split each sub-block into a local high-degree set (local core) and low-degree set (local periphery), and apply the same degree-preserving core-periphery swap restricted to edges whose endpoints lie inside that sub-block. Let $r_{\mathrm{cp}}^{(\ell)}$ be the per-node participation budget at level $\ell$, and let $\boldsymbol{\Omega}_{\mathrm{cp}}^{(\ell)}$ be the cumulative perturbation contributed by all swaps at that level.
Since every swap is a degree-preserving swap, the operator-norm control is identical at every level, by Proposition 2,
$\bigl\|\boldsymbol{\Omega}_{\mathrm{cp}}^{(\ell)}\bigr\|_{2}\;\leq\;2\,r_{\mathrm{cp}}^{(\ell)}$
After $L$ levels, the total perturbation is $\boldsymbol{\Omega}^{\mathrm{(tot)}}_{\mathrm{cp}}:=\sum_{\ell=0}^{L}\boldsymbol{\Omega}^{(\ell)}_{\mathrm{cp}}$, and each node participates in at most $r_{\mathrm{tot}}:=\sum_{\ell=0}^{L}r_{\mathrm{cp}}^{(\ell)}$ swaps overall, hence again (Proposition 2)
$\bigl\|\boldsymbol{\Omega}^{\mathrm{(tot)}}_{\mathrm{cp}}\bigr\|_{2}\;\leq\;2\,r^{\mathrm{tot}}_{cp}$
Therefore, whenever $2\kappa r_{\mathrm{tot}}/\gamma<1$, Proposition 3 yields the single-line eigenvector-rotation bound
$\sin\theta^{\mathrm{(tot)}}_{\mathrm{cp}}\;\leq\;\frac{2\kappa\,r^{\mathrm{tot}}_{\mathrm{cp}}}{\gamma}$
Because every swap preserves the in- and out-degree sequences, the degree proxy $\mathbf{d}$ is unchanged throughout, and in the neutral regime (Proposition 1) this implies that even a multi-level core-periphery construction keeps the leading eigenvector close to degree as long as the total per-node swap budget $r_{\mathrm{tot}}$ remains modest relative to $\gamma/\kappa$.
Remark 3 (Moment bound for the angle via core–periphery contrast).
Let $\phi_{\mathrm{cp}}(\mathbf{A};\mathcal{H},\mathcal{L})$ be the core–periphery contrast from Def 9, computed using the degree-based partition $(\mathcal{H},\mathcal{L})$. Along the degree-preserving core–periphery rewiring trajectory $\{\mathbf{A}^{(t)}\}_{t\geq 0}$ define
$\phi_{t}:=\phi_{\mathrm{cp}}(\mathbf{A}^{(t)};\mathcal{H},\mathcal{L}),\qquad\theta_{t}:=\theta(\mathbf{A}^{(t)})$
By construction, each accepted swap deletes an $\mathcal{L}\!\to\!\mathcal{L}$ edge and does not create a new $\mathcal{L}\!\to\!\mathcal{L}$ edge; thus $e_{\mathcal{L}\mathcal{L}}$ weakly decreases, and since $\phi_{\mathrm{cp}}(\cdot;\mathcal{H},\mathcal{L})$ is monotone in $e_{\mathcal{L}\mathcal{L}}$ when degrees are fixed, we have $\phi_{t+1}\geq\phi_{t}$ pathwise. If the affine upper comparison assumed in Proposition 5 holds on the reachable set for $\phi=\phi_{\mathrm{cp}}(\cdot;\mathcal{H},\mathcal{L})$, then Corollary 1 implies the one-moment control
$\mathbb{E}[\theta_{t}]\;\leq\;M_{\mathrm{cp}}\Bigl(1+\mathbb{E}\bigl[|\phi_{t}|\bigr]\Bigr)$
for a uniform constant $M_{\mathrm{cp}}$.
(Core–periphery-specific scaling of $M_{\mathrm{cp}}$.) Here the relevant budget is $r_{\mathrm{cp}}$ (as it enters the operator-norm control of the cumulative perturbation), and the partition enters only through $|\mathcal{H}|,|\mathcal{L}|$. A single accepted swap changes only $O(1)$ directed edges; in terms of the contrast statistic, the largest change comes from redirecting mass away from $\mathcal{L}\!\to\!\mathcal{L}$ and into blocks involving $\mathcal{H}$, which is controlled by the extremal degree product $d_{\max}^{\mathrm{out}}d_{\max}^{\mathrm{in}}$ (up to the normalization in Def 9). Thus the natural leverage scale is
$\Lambda_{\mathrm{cp}}\;\asymp\;C_{\mathrm{cp}}(\mathcal{H},\mathcal{L})\,d_{\max}^{\mathrm{out}}d_{\max}^{\mathrm{in}}$
with $C_{\mathrm{cp}}(\mathcal{H},\mathcal{L})$ collecting the block-size/normalization factors. Combining this with the perturbation factor $\kappa_{*}/\gamma_{*}$ and the budget $r_{\mathrm{cp}}$ gives the multiplicative envelope
$M_{\mathrm{cp}}\;\lesssim\;\widetilde{C}_{\mathrm{cp}}(\mathcal{H},\mathcal{L})\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\mathrm{cp}}\,\Lambda_{\mathrm{cp}}$
where $\widetilde{C}_{\mathrm{cp}}(\mathcal{H},\mathcal{L})$ is chosen large enough to absorb normalization conventions and any residual slack terms from the affine comparison.
If the in/out maxima are comparable and the common degree tail is power-law with exponent $\alpha>1$ (so $d_{\max}^{\mathrm{out}}\asymp d_{\max}^{\mathrm{in}}\asymp d_{\max}\asymp n^{1/\alpha}$ up to slowly varying factors), then $\Lambda_{\mathrm{cp}}\asymp C_{\mathrm{cp}}(\mathcal{H},\mathcal{L})\,d_{\max}^{2}$ and hence
$M_{\mathrm{cp}}\;\lesssim\;\widehat{C}_{\mathrm{cp}}(\mathcal{H},\mathcal{L})\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\mathrm{cp}}\,n^{2/\alpha}$
so heavier tails (smaller $\alpha$) enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ through increased extremal-degree leverage under degree-preserving swaps. $\circ$
5.3 Cycles
Definition 10 (Cycle-density statistic ($k$-cycle participation)).
A directed $k$-cycle is a simple motif $i_{1}\!\to i_{2}\!\to\cdots\to i_{k}\!\to i_{1}$ with distinct vertices. To measure a cycle, fix an integer $k\geq 3$ and let $\mathbf{A}\in\{0,1\}^{n\times n}$ be the adjacency matrix of a directed simple graph on vertex set $\{1,\dots,n\}$. A directed $k$-cycle is an ordered $k$-tuple of distinct vertices $(i_{1},\dots,i_{k})$ such that
$(i_{1}\!\to i_{2}),\ (i_{2}\!\to i_{3}),\ \dots,\ (i_{k-1}\!\to i_{k}),\ (i_{k}\!\to i_{1})\in\mathcal{E}(\mathbf{A})$
Let $C_{k}(\mathbf{A})$ denote the total number of directed $k$-cycles in $\mathbf{A}$ (counted up to cyclic rotation). [^15] Equivalently, $C_{k}(\mathbf{A})=\frac{1}{k}\,\#\{(i_{1},\dots,i_{k}):\ (i_{1},\dots,i_{k})\ \text{forms a directed $k$-cycle}\}$. Any consistent counting convention (e.g. counting ordered cycles and dividing by $k$) may be used. Define the $k$-cycle density by
$\phi_{k\text{-}\mathrm{cyc}}(\mathbf{A})\;:=\;\frac{C_{k}(\mathbf{A})}{(n)_{k}/k}\;\in\;[0,1],\qquad(n)_{k}:=n(n-1)\cdots(n-k+1)$
Here $(n)_{k}/k$ is the number of distinct directed $k$-cycles in the complete directed graph (with no self-loops), so $\phi_{k\text{-}\mathrm{cyc}}(\mathbf{A})$ is the fraction of all possible directed $k$-cycles that are present in $\mathbf{A}$. $\circ$
We now describe degree-preserving swaps that create short cycles (triangles and, more generally, $k$-cycles) from the neutral benchmark, and then invoke the same perturbation machinery as before: each cycle-forming move is a degree-preserving swap, hence contributes a rank-one $\{-1,0,1\}$ perturbation with spectral norm at most $2$. Naturally then the cumulative perturbation will be controlled by Proposition 2 and the resulting eigenvector rotation by the Stewart–Sun bound (Result 1).
Triangle-forming swap
Given $\mathbf{A}$, choose distinct $a,b,c$ with $(a\!\to b),(b\!\to c)\in\mathcal{E}(\mathbf{A})$ and $(c\!\to a)\notin\mathcal{E}(\mathbf{A})$. Pick edges $(c\!\to d)$ and $(e\!\to a)$ with $d,e\notin\{a,b,c\}$ such that $(c\!\to a)$ and $(e\!\to d)$ are admissible (no self-loops, not already present), and perform the swap
$(c\!\to d,\ e\!\to a)\quad\leadsto\quad(c\!\to a,\ e\!\to d)$
which preserves in- and out-degrees and closes the triangle $a\!\to b\!\to c\!\to a$. Writing $\boldsymbol{\Omega}^{(t)}_{\mathrm{tri}}$ for the cumulative perturbation after $t$ triangle-forming swaps and assuming each node participates in at most $r_{\mathrm{tri}}$ such swaps, Proposition 2 gives
$\bigl\|\boldsymbol{\Omega}^{(t)}_{\mathrm{tri}}\bigr\|_{2}\ \leq\ 2r_{\mathrm{tri}}$
Therefore, whenever $2r_{\mathrm{tri}}<\gamma/\kappa$, Result 1 and Proposition 3 yield the usual proxy-stability bound
$\sin\angle\!\bigl(\mathbf{d},\mathbf{v}^{(t)}\bigr)\ \lesssim\ \frac{2\kappa r_{\mathrm{tri}}}{\gamma}$
where $\mathbf{v}^{(t)}$ is the unit leading eigenvector of $\mathbf{A}^{(t)}$ and $\mathbf{d}$ is the (unit) degree proxy fixed by the degree sequence.
$\mathbf{k}$-cycles
Fix $k\geq 3$. To close an almost $k$-cycle $i_{1}\!\to\cdots\to i_{k}$ with missing edge $(i_{k}\!\to i_{1})$, perform the analogous swap that adds $(i_{k}\!\to i_{1})$ while deleting one outgoing edge of $i_{k}$ and one incoming edge of $i_{1}$ (and reconnecting their other endpoints). Each $k$-cycle-forming move is again a single degree-preserving swap, hence the same bounds hold with $r_{\mathrm{tri}}$ replaced by the per-node participation budget $r_{k\text{-}\mathrm{cyc}}$:
$\bigl\|\boldsymbol{\Omega}^{(t)}_{k\text{-}\mathrm{cyc}}\bigr\|_{2}\ \leq\ 2r_{k\text{-}\mathrm{cyc}},\qquad\sin\angle\!\bigl(\mathbf{d},\mathbf{v}^{(t)}\bigr)\ \lesssim\ \frac{2\kappa r_{k\text{-}\mathrm{cyc}}}{\gamma}$
whenever $2r_{k\text{-}\mathrm{cyc}}<\gamma/\kappa$.
Growing longer cycles
When ‘almost’ $k$-cycles are scarce, one can instead grow a triangle into a longer cycle by a constant number of additional swaps per unit increase in length (rerouting one cycle edge through a fresh directed length-$2$ chain disjoint from existing cycles). If creating a single length-$k$ cycle from a triangle uses at most $c_{k}=O(k)$ additional swapes, and each node participates in at most $r_{k}$ such growth steps, then the total perturbation satisfies the crude accumulation bound
$\bigl\|\boldsymbol{\Omega}^{\mathrm{tot}}_{k}\bigr\|_{2}\ \lesssim\ 2r_{\mathrm{tri}}+2c_{k}r_{k}$
and hence the Stewart–Sun angle control scales accordingly:
$\sin\angle\!\bigl(\mathbf{d},\mathbf{v}^{(k)}\bigr)\ \lesssim\ \frac{2\kappa}{\gamma}\,\bigl(r_{\mathrm{tri}}+c_{k}r_{k}\bigr)$
so longer cycles are ‘costlier’ because the number of required local swaps grows with $k$.
Remark 4 (Moment bound for the angle via the $k$-cycle statistic).
Fix $k\geq 3$ and let $\phi_{k\text{-}\mathrm{cyc}}(\mathbf{A})$ be the cycle statistic from Def 10. Along the $k$-cycle–forming, degree-preserving rewiring trajectory $\{\mathbf{A}^{(t)}\}_{t\geq 0}$ define
$\phi_{t}:=\phi_{k\text{-}\mathrm{cyc}}(\mathbf{A}^{(t)}),\qquad\theta_{t}:=\theta(\mathbf{A}^{(t)})$
The acceptance rule is $\phi_{k\text{-}\mathrm{cyc}}$-upper-contour (accepted swaps satisfy $\phi_{t+1}\geq\phi_{t}$), and by construction $0\leq\phi_{t}\leq 1$. If the affine upper comparison assumed in Proposition 5 holds on the reachable set for $\phi=\phi_{k\text{-}\mathrm{cyc}}$, then Corollary 1 yields
$\mathbb{E}[\theta_{t}]\;\leq\;M_{k\text{-}\mathrm{cyc}}\Bigl(1+\mathbb{E}\bigl[|\phi_{t}|\bigr]\Bigr)\;=\;M_{k\text{-}\mathrm{cyc}}\Bigl(1+\mathbb{E}[\phi_{t}]\Bigr)$
for a uniform constant $M_{k\text{-}\mathrm{cyc}}$.
($k$-cycle-specific scaling of $M_{k\text{-}\mathrm{cyc}}$.) The statistic depends only on the local closure of length-$k$ motifs, so a single accepted swap affects the cycle count through $O_{k}(1)$ edge incidences (a combinatorial constant depending only on $k$). Under degree preservation the extremal local edge-mass available for closing cycles is controlled by $d_{\max}^{\mathrm{out}}$ and $d_{\max}^{\mathrm{in}}$; in worst case, closing short cycles concentrates around high-degree vertices, giving the leverage scale
$\Lambda_{k\text{-}\mathrm{cyc}}\;\asymp\;C_{k}\,d_{\max}^{\mathrm{out}}d_{\max}^{\mathrm{in}},$
where $C_{k}$ absorbs the $k$-dependent combinatorial factor and the normalization convention used in $\phi_{k\text{-}\mathrm{cyc}}$. Combining this with the perturbation factor $\kappa_{*}/\gamma_{*}$ and the $k$-cycle participation/budget parameter $r_{k\text{-}\mathrm{cyc}}$ gives the multiplicative envelope
$M_{k\text{-}\mathrm{cyc}}\;\lesssim\;\widetilde{C}_{k}\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{k\text{-}\mathrm{cyc}}\,\Lambda_{k\text{-}\mathrm{cyc}}$
with $\widetilde{C}_{k}$ chosen large enough to absorb normalization and any residual slack from the affine comparison.
If the in/out maxima are comparable and the common degree tail is power-law with exponent $\alpha>1$ (so $d_{\max}^{\mathrm{out}}\asymp d_{\max}^{\mathrm{in}}\asymp d_{\max}\asymp n^{1/\alpha}$ up to slowly varying factors), then $\Lambda_{k\text{-}\mathrm{cyc}}\asymp C_{k}d_{\max}^{2}$ and hence
$M_{k\text{-}\mathrm{cyc}}\;\lesssim\;\widehat{C}_{k}\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{k\text{-}\mathrm{cyc}}\,n^{2/\alpha}$
so heavier tails (smaller $\alpha$) enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ by increasing the combinatorial leverage available for closing cycles around hubs. $\circ$
6 Heavy-tailed degrees and eigenvector localization
A subtle but practically important point is that the constants in our moment bounds are not universal, and this is exactly where the four statistics truly part ways. All four bounds share the same backbone: (i) eigenvector responsiveness enters through distortion and spectral-gap parameters $(\kappa,\gamma)$, and (ii) the cumulative size of the perturbation is controlled by the bounded-participation operator norm $r$. Where they differ is in the statistic-to-angle sensitivity—how changes in the chosen global statistic $\phi$ translate into bounds on the deviation angle $\theta(\mathbf{A})$. The algebraic form of the moment transfer is stable, but its scale is dictated by how sensitive a statistic is under degree-preserving swaps and, conversely, how much eigenvector rotation those swaps can actually induce. A convenient proxy for this sensitivity is given by the affine parameters, whose exact functional form and magnitude is inherently statistic-specific.
Put differently, there are really two layers at work. The first layer is purely spectral: $\kappa/\gamma$ measures how fragile the leading eigenvector is to perturbations of a given operator norm. The second layer is combinatorial/geometric: it tells you how much operator-norm perturbation (and hence how many degree-preserving edits, under the $r$-bounded participation rule) is needed to achieve a prescribed change in $\phi$. The statistics differ primarily in this second layer—some can be moved substantially by rearranging a small number of edges concentrated around a few vertices, while others only shift appreciably after many more, and more dispersed, changes.
Assortativity.
For assortativity, the affine constants in our moment bounds are pinned down by endpoint-degree dispersion. Newman assortativity is a normalized covariance of degrees across edges, and $\phi_{p,q}$ simply correlates the $p$-degree at the tail of an edge with the $q$-degree at its head (out/in, in/in, etc.). A degree-preserving swap therefore cannot change the endpoint-degree multisets. It can only change ‘who is paired with whom’ across edges, i.e. the cross-edge covariance itself. Which is exactly what assortativity records.
The normalization $\nu_{p,q}=1/(\sigma_{T}\sigma_{H})$ turns the raw degree-product covariance into a dimensionless correlation, where $\sigma_{T}$ and $\sigma_{H}$ are the standard deviations of the tail- and head-side endpoint-degree samples. Intuitively, this fixes the ‘units’ of mixing: when endpoint degrees are highly dispersed, the same absolute change in the degree-product sum produces a smaller standardized change in $\phi_{p,q}$. Along our rewiring trajectory, $\sigma_{T}$ and $\sigma_{H}$ (hence $\nu_{p,q}$) remain invariant, so the scale is set once and for all by the degree distribution, while rewiring acts only through changes in the cross-edge covariance.
Finally, because each swap re-pairs only a constant number of endpoints, the per-swap movement in $\phi_{p,q}$ is driven by the upper tail: the largest increments come from reassigning edges incident to rare high-degree vertices, which generate the biggest changes in the degree-product sum. Put together, this is why assortativity is usually slow to move: each swap re-pairs only a few endpoints, and the statistic averages over all edges, so a visible change typically requires many coordinated re-pairings. In heavy-tailed networks this picture becomes more uneven. High-degree endpoints are rare, so most swaps touch only moderate degrees and produce tiny changes. But when a swap does involve a rare high-degree node, the degree–product sum can jump, and so can $\phi_{p,q}$. In particular, heavier tails tend to worsen the worst-case affine constants in our moment bounds: even though greater dispersion reduces the standardized effect of a typical swap (via $\nu_{p,q}=1/(\sigma_{T}\sigma_{H})$), the presence of more extreme high-degree endpoints increases the size of the occasional jumps that drive the upper envelope. Therefore, in the assortativity case, the affine constants that bound the first two moments are controlled by endpoint-degree dispersion (through $\sigma_{T}\sigma_{H}$ and $\nu_{p,q}$) together with how much mass sits in the upper tail, since these two ingredients determine both the typical increment size and the occasional large jumps by which degree-preserving swaps can move $\phi_{p,q}$.
Communities.
For community structure, the affine constants are shaped by the chosen partition because $\phi_{\mathrm{com}}$ is a mass-transfer statistic, i.e. it measures how much edge mass lies within blocks relative to across blocks. A degree-preserving swap cannot change how many stubs each vertex contributes, so the only way to increase $\phi_{\mathrm{com}}$ is to re-route existing cross-block edges into within-block positions while keeping every vertex’s degree fixed. In that sense the rewiring is a constrained transport problem: one is trying to ‘pour’ as much mass as possible into the diagonal blocks of the adjacency matrix without changing the row/column sums.
How fast this transfer can proceed depends on two interacting ingredients. First, the partition geometry—in particular the number of communities and the resulting block sizes—controls the combinatorial supply of admissible within-block endpoints. With many communities (hence smaller blocks), within-block room is scarcer, so the process saturates sooner. With fewer, larger (or more balanced) blocks, there are simply more eligible within-block pairs to absorb diverted mass. Second, the degree tail matters because high-degree vertices dominate the supply and demand of edge mass: most of what can be moved is carried by the few large-degree nodes. If these hubs are distributed across blocks in a way that gives them abundant within-block partners, cross-block mass can be converted quickly. If hub degree mass is concentrated in blocks with limited within-block capacity, gains in $\phi_{\mathrm{com}}$ stall early even if many swaps remain available.
This is why, in the community case, the bounding affine constants depend on the interaction between the number of communities (which sets within-block capacity) and the tail of the degree distribution (which concentrates the movable mass): the former determines how much within-block ‘space’ exists, and the latter determines how much mass is trying to flow into it.
Core-periphery.
Core-periphery is similar in spirit to communities in that we relocate edge mass under degree preservation, but the target is asymmetric. The statistic $\phi_{\mathrm{cp}}$ rises mainly by suppressing periphery–periphery adjacency: we drain the $\mathcal{L}\!\to\!\mathcal{L}$ block and redirect that mass so that periphery endpoints connect through the core $\mathcal{H}$. Since out- and in-stub counts are fixed vertexwise, rewiring cannot change how much periphery stub mass exists. It can only change where it lands. Increasing $\phi_{\mathrm{cp}}$ is therefore a progressive transfer of $\mathcal{L}\!\to\!\mathcal{L}$ edges into $\mathcal{L}\!\to\!\mathcal{H}$, $\mathcal{H}\!\to\!\mathcal{L}$, and/or $\mathcal{H}\!\to\!\mathcal{H}$ locations.
The affine bounding constants are governed by the split $(\mathcal{H},\mathcal{L})$ and by how long this drainage can continue before capacity constraints bite. Under degree preservation, the core has a fixed aggregate ability to absorb redirected periphery edge-ends. If that capacity is insufficient, a nontrivial residue of $\mathcal{L}\!\to\!\mathcal{L}$ mass is unavoidable. Degree dispersion matters here too: in heavy-tailed networks a few very high-degree core vertices act as ‘large-capacity hubs’, creating more room to reattach periphery edge-ends to $\mathcal{H}$ before saturation.
The core-periphery statistic can have high leverage: a modest number of well-chosen swaps can rapidly funnel walk mass through $\mathcal{H}$, a structural asymmetry to which the Perron vector is particularly responsive. Which is why the affine bounds in the core-periphery case depend on the interaction between $(\mathcal{H},\mathcal{L})$ split and the heaviness of the degree tail, since the two together determine how much periphery mass can be redirected through the core and how sharply paths can be concentrated therein.
Cycles.
For $k$-cycles, $\phi_{k\text{-}\mathrm{cyc}}$ is motif-based and the relevant rewiring is inherently local: a single accepted swap can create a new short cycle (or extend an incipient one) by rearranging only a constant number of edges, without any global reorganization of mixing patterns. This locality is precisely why cycles are a useful stress-test for eigenvector robustness: inserting a small motif in a small region can induce a noticeable, localized eigenvector distortion even when coarse global summaries move slowly. Degree dispersion still matters at this combinatorial level: high-degree vertices offer a large menu of neighbors, so when a swap touches a hub it is often easier to ‘close’ a cycle around it simply because there are many more admissible ways to complete the required adjacency pattern under degree preservation.
Longer cycles are typically costlier because a $k$-cycle requires coordinating adjacency among $O(k)$ vertices. In practice one often has to stage the motif over multiple accepted swaps, so the per-swap increase in $\phi_{k\text{-}\mathrm{cyc}}$ tends to shrink as $k$ grows. Heavy tails, however, reshape this cost: hubs supply abundant ‘attachment points’, so long cycles can be stitched together with far less coordination than in thin-tailed graphs. In this sense, the longer the target cycle, the more the tail thickness matters as hubs are precisely what make many kinds of long-cycle constructions feasible.
6.1 Heavy tails and the moment bounds
A unifying theme across all four statistics is that the heaviness of the degree tail governs how much structural change one can generate with a fixed degree-preserving perturbation budget. When the degree distribution is heavy-tailed, a small set of hubs carries a disproportionate share of edge-ends. This concentration creates two generic effects that recur throughout our constructions. First, it generates high leverage: swaps that involve hubs can move the statistic substantially, because much of the relevant mass transfer, edge pairing, or motif closure is mediated by these few vertices. Second, it makes the evolution lumpy: most swaps that touch moderate-degree nodes do little, but the occasional hub-involving swap can produce a visible jump. In worst-case terms, this is precisely what inflates the affine constants—heavier tails expand the range of one-step (and short-horizon) changes achievable under the same degree-preserving operator-norm budget, and hence widen the envelope of possible eigenvector rotations.
This lens helps reconcile the four cases. For assortativity, heavy tails amplify the rare but large covariance jumps created by re-pairing edges incident to hubs (even as tail dispersion also affects the normalization). For communities and core–periphery, heavy tails place much of the ‘movable mass’ in a few vertices, so the speed of mass transfer is governed by whether the partition/core split gives those hubs enough admissible destinations before capacity saturates. For cycles, tail heaviness matters combinatorially: hubs offer many local routes, making it easier not only to close short cycles but also to assemble and stitch together longer cycles, since hubs supply abundant attachment points for coordinating adjacency across many vertices.
There is, however, a further (and subtler) channel through which tail heaviness can enter the moment bounds: it may also affect the spectral sensitivity parameters themselves. Hub-centered rewiring typically builds multiple forms of structure around the same small set of vertices (with many low-degree neighbors acting as ‘feeders’ or attachment points), and this can make distinct eigenmodes less separable. In the language of the Stewart–Sun bound, this corresponds to potential increases in the distortion factor $\kappa$ (a more ill-conditioned eigenbasis) and/or decreases in the spectral gap $\gamma$ (weaker separation between $\lambda_{1}$ and the rest). Intuitively, when a low-degree vertex is embedded into a hub-centered local structure, its influence is no longer well summarized by its degree alone: it can ‘borrow’ the hub’s amplification along walks, so perturbations supported on such mixed hub–periphery motifs are more easily transmitted across modes. The resulting modal interaction is exactly the regime in which $\kappa$ can grow and $\gamma$ can shrink [^16] In general $\kappa(A)$ and $\gamma(A)$ are not invariant under degree-preserving rewiring: they depend on the current adjacency matrix and may drift along the trajectory. Accordingly, as in earlier remarks, we interpret the bound uniformly on a reachable region $S_{0}$, working with time-invariant descriptors $\kappa^{\ast}:=\sup_{A\in S_{0}}\kappa(A)$ and $\gamma^{\ast}:=\inf_{A\in S_{0}}\gamma(A)>0$. In heavy-tailed settings these suprema/infima can be larger (worse), since hub-centered local-structure formation can increase modal interaction (raising $\kappa(A)$ and potentially shrinking $\gamma(A)$)..
These considerations point to two opposing forces. The direct effect of heavy tails is leverage: for a fixed perturbation budget they typically generate larger changes in the statistic—equivalently, for a fixed change in the statistic they can require fewer accepted swaps and hence permit smaller angle deviation. The indirect effect is spectral: if hub-centered structures worsen conditioning (larger $\kappa$ and/or smaller $\gamma$), then a given perturbation can translate into a larger eigenvector rotation—equivalently, a fixed change in the statistic can be accompanied by a larger angle deviation. The net effect of degree-tail heaviness on the angle–moment bounds is therefore a priori ambiguous, and depends on which channel dominates in the regime of interest.
7 Concluding remarks
Applied researchers working in data-scarce environments seldom have the information necessary to compute the eigenvectors of the network. Since the eigenvector is a good description of the systemic significance of nodes—whether it be firms about to foreclose or humans about to transmit a disease—the researcher is left with having to proxy it. Which he typically does with the degree vector [^17] Typically, the number of connections of a node is more easily known than the entire network structure.. Over the years, a certain ‘folk theorem’ has emerged around this substitution, which claims that the procedure is not all too bad when the network has neutral degree-mixing and lacks meso structures. This folk theorem, however, does not tell us anything about the error introduced by the substitution in networks that violate the neutrality assumptions. This paper presents analytical bounds on the error that is born from substituting the degree-vector for the eigenvector in assortative networks with a modicum of meso structures. We start from a neutral benchmark in which degree and eigenvector centrality align, and then introduce four common departures from neutrality: degree assortativity, community structure, core-periphery organization, and short directed cycles. For each departure, we describe a degree-preserving rewiring mechanism that strengthens the corresponding structure, construct the induced perturbation of the adjacency matrix, and translate it into an upper bound on the deviation angle between the degree vector and the leading eigenvector via the Stewart-Sun Perturbation Bound. With this procedure, we placed a boundary around an error term of some empirical importance in economics, epidemiology, opinion dynamics, and other applied sciences.
Note that our analytical results are only reasonable upper bounds. In principle, local structures can be introduced without disturbing the alignment between degree and eigenvector. In fact, one could even place local structures to bring the two closer. The reason is simple: insofar as the preponderance of local structure accrues to (or is concentrated around) high-degree nodes, its presence need not pull the Perron eigenvector away from the degree proxy. One reason for such an aligned preponderance is the combinatorial fact that higher-degree nodes have a greater possibility of participating in a wide variety of local structures. In fact, some local structures—such as long cycles—are difficult to construct without the participation of high-degree nodes. So at one extreme, local structures are no trouble at all. At the other extreme, they can wholly misalign the degree and the eigenvector. To study this misalignment, we ensure that the local structures created through our rewiring process maximally disturb the alignment between degree and eigenvector. We do this through two assumptions. The first of which is that no node can participate in more than $r$ rewiring with $r\ll n$ [^18] This assumption is required for the Stewart-Sun perturbation bound to hold.. Naturally, this means that high-degree nodes cannot participate in as many local structures as they might have otherwise. The limit curtails the number of local structures that can form around high-degree nodes, thereby causing greater deviation of the eigenvector from the degree vector. The bound on the angle of deviation derived using the Stewart-Sun procedure is therefore naturally an envelope, particularly in the case of heavy-tailed networks where the limit $r$ will almost certainly be binding. The second assumption involves putting structure on the evolution of the angle of deviation in response to the evolution of statistics that measure assortativity and local structures. We assume that the expectation of the angle of deviation increases with the concerned network statistics. This structural assumption allows bound the moments of the angle of deviation with the moments of the concerned network statistics. Naturally, this expectational assumption means that this moment-bound, too, is an envelope. All of this is to say that our procedure establishes a maximal angle of deviation between the degree vector and the eigenvector, a ‘reasonable’ worst-case scenario if you will [^19] We say ‘reasonable’ worst-case because one could develop even ‘worse’ worse-case scenarios by implanting local structures at particular chosen locations in a graph.. In real-world networks, the error induced by substituting the degree vector with the eigenvector could be markedly smaller.